# and quantum information science

### v2

@@ 7.2 Time-dependent properties of the holographic proposals
One may worry that this quantity is divergent, but it turns out that its time dependent part $\langle\ell\rangle-\langle\ell\rangle_{t=0}$, which is all we need, is finite and independent of the regulator. Denoting by $S_{0}$ the leading-order entropy of a black hole in JT gravity, the ERB’s length grows linearly until it saturates at a time and value both of order $e^{S_{0}}$ (see Fig. 1 and Fig. 5 of [31]). This behavior matches the time evolution of quantum complexity depicted in Fig. 3, but with a difference. The variance of the ERB’s length is negligeble at times $t\sim\mathcal{O}(e^{S_{0}})$, but it monotonically grows until becoming of the same order as $\langle\ell\rangle$ at times $t\sim\mathcal{O}(e^{2S_{0}})$.
Recently, the authors of [269] made further progress in defining a nonperturbative version of holographic complexity. Rather than focusing on the length expectation value (124), they studied the spectral decomposition of $\langle e^{-\Delta\ell}\rangle$, and identified the latter quantity as a generating function to calculate quantum complexity. The generating function shows a slope-ramp-plateau structure similar to the case of the SFF, see Figs. 6-7 of [269]. In the limit $\Delta\rightarrow 0$, the ramp disappears from the time evolution (Fig. 12 of [269]), leading to the characteristic linear growth and late time saturation of complexity. Additional definitions leading to a late time saturation are based on spread complexity [35], and will be discussed in Sec. 7.5.1.
A general mechanism for saturation of Einstein-Rosen Bridges (ERBs) in any dimension was proposed in [36, 37], the authors of which constructed a basis of microstates for eternal black holes in any dimension, and for any theory described at low energies by general relativity. These states, consisting of ERBs of different lengths supported internally by shells of matter, are perturbatively orthogonal. Hence, they naively describe an infinite space of ERBs of arbitrary lengths. However, topology-changing wormholes in the gravitational path integral lead to small quantum overlaps which make these states linearly dependent, so that they span a Hilbert space of dimension precisely equal to the exponential of the Bekenstein-Hawking entropy. This means that we can pick a complete basis of wormholes of bounded length; so, long wormholes of the kind that appear at late times in the classical geometry can be regarded as superpositions of short bridges. This suggests a quantum mechanical saturation of wormhole length in any dimension, similarly to the saturation mechanism for the thermal two-point function in 2d JT gravity via tunnelling to baby universes with shorter ERBs [270]. If long ERBs should be regarded as superpositions of small bridges in this way, we should also ask if there is any definite notion of wormhole volume that can be identified directly with complexity. We will answer this question in Sec. 7.5.1.
@@ 7.3 Time-independent properties of the holographic proposals
A similar divergent behavior appears when studying the entanglement entropy. In that case, the leading divergence behaves as an area law [^30] Note that area law here refers to the term dominant in a short distance UV cutoff expansion (small $\delta$), not the term dominant in a large entangling region volume limit, as is sometimes the case in the condensed matter literature. The two expansions can differ when an additional scale is present, e.g., a temperature. $A/\delta^{d-2}$ with $A$ the area of the entangling surface, $\delta$ a short distance cutoff and $d$ the spacetime dimension of the boundary field theory. This is true both in holographic settings and in free field theories [271, 272]. Furthermore, in the case of a smooth entangling surface, the divergences jump in powers of two, capturing different geometric features of the entangling surface. Universal contributions, depending on the anomalies of the theory, appear as coefficients of either the logarithmic or the constant parts of the expansion in the short-distance cutoff.
The above results motivated the study of UV divergences for the complexity of a vacuum state. In [273], the authors explored the CV and CA proposals in empty AdS spacetime, see Fig. 12. They found that these quantities always diverge proportionally to the volume $\mathcal{V}$ of the system as
@@
Figure 12: Illustration of the CV (left) and CA (right) proposals in empty AdS spacetime for a state living on a constant-time boundary slice. The geometric quantities are evaluated inside the brown region with ends fixed by the short distance cutoff, illustrated in the figure as the pink region. Figure adapted from [274].
Next, one can consider complexity for more general states, such as the thermofield-double state dual to the two-sided eternal black hole (119). In this case, one defines the complexity of formation as the excess complexity required to construct this state compared to the complexity of two copies of the vacuum, i.e., $\mathcal{C}_{\text{form}}\coloneqq\mathcal{C}(|\mathrm{TFD}(t=0)\rangle)-2\mathcal{C}(|\text{vac}\rangle)$. This quantity was evaluated in [274] using the CV and CA proposals and it was found that in the high temperature regime and in dimensions $d>2$, it grows linearly with the entropy of the system, i.e., $\delta\mathcal{C}=k_{d}S$. Conversely, when $d=2$ the complexity was found to be a fixed constant, independent of the temperature.
@@ 7.4 Generalizations of the holographic proposals
We focus on three generalizations: (1) the extension to the case of subregions, (2) settings with defects or boundaries, and (3) applications to de Sitter spacetime. It is harder to define and test ideas about holographic complexity in these settings than in the standard ones discussed in the previous subsections. Nevertheless, they add information about which notion of quantum complexity should be expected on the field theory side of the holographic duality. [^31] Field theory studies of complexity related to these three extensions appear in, e.g., [275, 110] for subregions, and [276] for defects. See also [277] for a relevant tensor network construction and [278] for quantum duals to de Sitter spacetime. These constructions usually involve simplified models such as counting arguments for random circuits or even Gaussian free field theories and so can only be compared to the holographic results in a qualitative way.
@@
Consider a mixed state associated with a subregion $\mathcal{A}$ of the boundary. The CV proposal can be extended to this case by computing the maximal volume of a codimension-one surface delimited by the subregion $\mathcal{A}$ and its HRT surface [279]. According to the bulk reconstruction paradigm (see [280] for a review), the information carried by the reduced density matrix $\rho_{\mathcal{A}}$ of a mixed state is holographically encoded by the entanglement wedge (EW), which is the bulk domain of dependence of a region that is enclosed within $\mathcal{A}$ and its HRT surface [281]. Therefore, in the case of CA (CV2.0) conjecture, subregion complexity was defined as the gravitational action (spacetime volume) in the bulk region given by the intersection of the WDW patch with the Entanglement Wedge [252].
@@
The general structure of the CV and CA proposals has been classified for a ball-shaped subregion $\mathcal{B}$ in arbitrary dimensions, giving a leading UV divergence similar to the full system [252]. That is, $\mathcal{C}_{\rm sub}\propto k_{d}C_{T}\mathcal{V}(\mathcal{B})/\delta^{d-1},$ where $k_{d}$, the same constant as in eq. (125), carries information on various ambiguities and the dimensionality of the system, $C_{T}$ is the central charge, and $\mathcal{V}(\mathcal{B})$ is the volume of the ball. In general, the subregion complexity proposals have been shown to be superadditive in a pure state; [^32] Consider any holographic complexity conjecture $\mathcal{C}(\Sigma)$ defined on a spacetime region anchored to a boundary Cauchy slice $\Sigma$. Let us split the Cauchy slice $\Sigma$ into a subregion $\mathcal{A}$ and its complement $\bar{\mathcal{A}}$. Denote with $\mathcal{C}(\mathcal{A})$ and $\mathcal{C}(\bar{\mathcal{A}})$ the corresponding subregion complexities, we say that holographic complexity is superadditive if $\mathcal{C}(\Sigma)>\mathcal{C}(\mathcal{A})+\mathcal{C}(\bar{\mathcal{A}})\,,$ (126) and subadditive if the inequality is inverted. CA is an exception, since one can tune the counter term scale $L_{\rm{ct}}$ for the null-boundaries (defined below equation (114)) to either get a subadditive or superadditive behaviour [110, 275, 282, 283, 284]. However, restricting the counter term scale $L_{\rm{ct}}$ such that the leading divergence in CA is positive, results in CA being superadditive too. Subsequent studies of holographic complexity associated with subregions were carried out, e.g., in [279, 252, 285, 286, 287, 288, 289, 275, 282, 290, 291, 292, 293, 110, 284, 294, 295, 296].
@@
The proposed formulae for subregion complexity have an interesting behavior in the BTZ black hole geometry. Given a boundary subregion composed by $q$ segments of total length $\ell_{\rm tot}$, the CV proposal reads [286]
@@
where $c$ is the central charge, $\chi$ the Euler characteristic of the maximal surface, and $\delta$ a UV cutoff regulator. Interestingly, this formula is topological, i.e., it does not depend on the temperature. This feature is not shared by subregion CV in higher dimensions [279, 285].
The subregion CA and CV2.0 conjectures for a single segment of length $\ell$ in the BTZ background were studied in [294]
@@
where $L_{\rm ct}$ is a length scale introduced by a counterterm on null surfaces in the gravitational action, see the discussion below Eq. (114), and $S_{\rm EE}$ is the entanglement entropy of the segment. These results present an appealing feature in that the complexity contains a piece proportional to the entanglement entropy plus an additional constant which could naively depend on the topology as in the CV conjecture results (127). However, an analysis of CA and CV2.0 conjectures in the case of multiple segments in the BTZ background reveals an intricate dependence on the subregion sizes and temperature (see Eqs. (4.23)–(4.24) of [294]), which does not follow the same pattern for the dependence of complexity on the entanglement entropies as the one in Eqs. (128a)–(128b) [294, 126]. Moreover, we do not expect a simple linear relation between complexity and entanglement entropy in dynamical situations, since, as we argued at the beginning of Sec. 7 the two quantities evolve very differently with time.
@@
The first is the two-sided Randall-Sundrum (RS) model, a gravity solution where two patches of AdS spacetime are glued together along a thin codimension-one brane [276]. The second is the AdS/BCFT model, where an end-of-the-world brane, corresponding to a boundary in the dual CFT, delimits vacuum AdS [297, 298]. The last is Janus AdS, a dilatonic deformation of AdS spacetime dual to an interface CFT and arising from the dimensional reduction of a solution in type IIB supergravity [299, 300].
@@
$2$-sided Randall-Sundrum $\frac{2}{3}c\,\eta_{\rm RS}\,\log\left(\frac{\ell}{\,\delta}\right)+\text{finite}$ 00
@@
Table 1: Subregion complexity of formation $\Delta\mathcal{C}$ for an interval of length $\ell$ in the cases of CV and CA, compared to empty AdS spacetime. The coefficients of the log divergences $\eta,\tilde{\eta}$ depend on the details of the defect or boundary, see [276, 297, 298, 299, 300].
Holographic complexity was studied in Janus AdS geometry under different regularization schemes [301]. In dimensions $d=2,4$, the authors found a scheme-independent logarithmic divergence, while in odd-dimensions a scheme-independent finite term was observed instead. This result parallels the behavior of the entanglement entropy of smooth subregions, see e.g., [272].
@@
De Sitter spacetime is a maximally symmetric solution of Einstein’s equations with a positive cosmological constant. It describes the early stages of evolution of our universe, and possibly the late stages if there is a positive cosmological constant (e.g., see the reviews [302, 303, 304]). Since there is no timelike boundary in de Sitter spacetime, it is challenging to find a holographic interpretation of this geometry in terms of a microscopic quantum mechanical system. Nevertheless, the physical relevance of this spacetime stimulated several attempts [305, 306, 307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 327, 328, 329, 330, 331, 332, 192, 333, 334, 335, 336, 337, 338, 339, 278, 340, 341, 342, 343, 344, 345, 346, 347, 348, 349, 350]. Recent developments suggest that the inclusion of an artificial timelike boundary in de Sitter spacetime is needed for the consistency of thermodynamics [340], microstate counting [330, 336, 347], and to take into account the role played by an observer [324, 342, 343, 344, 351, 352, 350]. These observations led to a revival of static patch holography, according to which a dual quantum theory lives on a codimension-one timelike surface (the stretched horizon) located just inside the cosmological horizon [308, 319, 322, 329, 338, 337, 335, 339, 345, 346, 348, 349]. The holographic complexity proposals introduced in subsection 7.1 can be applied to de Sitter spacetime by requiring that the geometric objects of interest are anchored at the stretched horizons on the two sides of the geometry [337]. The study of these generalizations is speculative, but may hint towards the expected properties of the quantum mechanical dual to de Sitter space.
@@
A main novelty, compared to the AdS case, is that CV, CV2.0 and CA conjectures in dS space include contributions from spacetime regions close to timelike infinity, and therefore diverge at a finite critical time, e.g., see Eq. (3.17) of [353]. [^33] One can regularize both holographic complexity and its rate by introducing a cutoff surface at future infinity of dS spacetime. After introducing the cutoff, holographic complexity grows linearly, e.g., see Eq. (3.30) of [353]. It was argued in [337] that this behavior should be interpreted as a hyperfast growth of complexity corresponding to circuits where the gates involve a large number of qubits at each step in the time evolution (e.g., see [278]). The hyperfast growth also occurs in two dimensions [354] and in models of inflation where a bubble of de Sitter spacetime is inside an AdS geometry [355]. [^34] Note however, that in two dimensions, the CV proposal yields a diverging rate of change in complexity as the critical time is approached, but the complexity itself remains finite in this limit. [^35] Holographic complexity has also been studied in other cosmological models, for instance the case of Kasner-like singularities [356]. However, there are two significant exceptions to this trend. First, there is a class of codimension-one CAny observables that exhibits linear or exponential growth persisting forever, without a divergent rate at finite time [357]. Furthermore, in any dimension of empty de Sitter space, the volume observable relevant to the CV conjecture does not strictly speaking “exist” if we anchor the relevant surfaces on the observer’s world line. To address this, we can instead anchor the surface to a “stretched horizon” at some radius $r_{\text{st}}$. But in this case the volume stops “existing” at a time $t_{\rm crit}$ proportional to $L\,\mathrm{arctanh}(r_{\text{st}}/L)$, where $L$ is the de Sitter scale. So, in addition to the dependence on $L$ and therefore on temperature, the critical time also depends on the radius of the “stretched horizon” that regularizes the problem. If the stretched horizon is very close to the de Sitter horizon, it can be made arbitrarily large [353]. Similarly, the CV conjecture is strictly speaking only defined for a short time of the order of the inverse temperature in the case of two-dimensional centaur geometries, i.e., gravitational models where de Sitter spacetime is glued to an asymptotic AdS region with a standard timelike boundary [358], because the maximal volume surfaces no longer exist after the critical time unless they are regularized in some way. It would be interesting to understand whether these behaviors of the CV observables have an interpretable origin in a dual description.

### v3

@@ 7.2 Time-dependent properties of the holographic proposals
One may worry that this quantity is divergent, but it turns out that its time dependent part $\langle\ell\rangle-\langle\ell\rangle_{t=0}$, which is all we need, is finite and independent of the regulator. Denoting by $S_{0}$ the leading-order entropy of a black hole in JT gravity, the ERB’s length grows linearly until it saturates at a time and value both of order $e^{S_{0}}$ (see Fig. 1 and Fig. 5 of Iliesiu:2021ari). This behavior matches the time evolution of quantum complexity depicted in Fig. 3, but with a difference. The variance of the ERB’s length is negligeble at times $t\sim\mathcal{O}(e^{S_{0}})$, but it monotonically grows until becoming of the same order as $\langle\ell\rangle$ at times $t\sim\mathcal{O}(e^{2S_{0}})$.
Recently, the authors of Miyaji:2025yvm made further progress in defining a nonperturbative version of holographic complexity. Rather than focusing on the length expectation value (124), they studied the spectral decomposition of $\langle e^{-\Delta\ell}\rangle$, and identified the latter quantity as a generating function to calculate quantum complexity. The generating function shows a slope-ramp-plateau structure similar to the case of the SFF, see Figs. 6-7 of Miyaji:2025yvm. In the limit $\Delta\rightarrow 0$, the ramp disappears from the time evolution (Fig. 12 of Miyaji:2025yvm), leading to the characteristic linear growth and late time saturation of complexity. Additional definitions leading to a late time saturation are based on spread complexity Balasubramanian:2024lqk, and will be discussed in Sec. 7.5.1.
A general mechanism for saturation of Einstein-Rosen Bridges (ERBs) in any dimension was proposed in Balasubramanian:2022gmo, Balasubramanian:2022lnw, the authors of which constructed a basis of microstates for eternal black holes in any dimension, and for any theory described at low energies by general relativity. These states, consisting of ERBs of different lengths supported internally by shells of matter, are perturbatively orthogonal. Hence, they naively describe an infinite space of ERBs of arbitrary lengths. However, topology-changing wormholes in the gravitational path integral lead to small quantum overlaps which make these states linearly dependent, so that they span a Hilbert space of dimension precisely equal to the exponential of the Bekenstein-Hawking entropy. This means that we can pick a complete basis of wormholes of bounded length; so, long wormholes of the kind that appear at late times in the classical geometry can be regarded as superpositions of short bridges. This suggests a quantum mechanical saturation of wormhole length in any dimension, similarly to the saturation mechanism for the thermal two-point function in 2d JT gravity via tunnelling to baby universes with shorter ERBs Saad:2019pqd. If long ERBs should be regarded as superpositions of small bridges in this way, we should also ask if there is any definite notion of wormhole volume that can be identified directly with complexity. We will answer this question in Sec. 7.5.1.
@@ 7.3 Time-independent properties of the holographic proposals
A similar divergent behavior appears when studying the entanglement entropy. In that case, the leading divergence behaves as an area law [^33] Note that area law here refers to the term dominant in a short distance UV cutoff expansion (small $\delta$), not the term dominant in a large entangling region volume limit, as is sometimes the case in the condensed matter literature. The two expansions can differ when an additional scale is present, e.g., a temperature. $A/\delta^{d-2}$ with $A$ the area of the entangling surface, $\delta$ a short distance cutoff and $d$ the spacetime dimension of the boundary field theory. This is true both in holographic settings and in free field theories Nishioka2009RTrev, Casini2009RevFreeEnt. Furthermore, in the case of a smooth entangling surface, the divergences jump in powers of two, capturing different geometric features of the entangling surface. Universal contributions, depending on the anomalies of the theory, appear as coefficients of either the logarithmic or the constant parts of the expansion in the short-distance cutoff.
The above results motivated the study of UV divergences for the complexity of a vacuum state. In Carmi2017DivC, the authors explored the CV and CA proposals in empty AdS spacetime, see Fig. 12. They found that these quantities always diverge proportionally to the volume $\mathcal{V}$ of the system as
@@
Figure 12: Illustration of the CV (left) and CA (right) proposals in empty AdS spacetime for a state living on a constant-time boundary slice. The geometric quantities are evaluated inside the brown region with ends fixed by the short distance cutoff, illustrated in the figure as the pink region. Figure adapted from Chapman:2016hwi.
Next, one can consider complexity for more general states, such as the thermofield-double state dual to the two-sided eternal black hole (119). In this case, one defines the complexity of formation as the excess complexity required to construct this state compared to the complexity of two copies of the vacuum, i.e., $\mathcal{C}_{\text{form}}\coloneqq\mathcal{C}(|\mathrm{TFD}(t=0)\rangle)-2\mathcal{C}(|\text{vac}\rangle)$. This quantity was evaluated in Chapman:2016hwi using the CV and CA proposals and it was found that in the high temperature regime and in dimensions $d>2$, it grows linearly with the entropy of the system, i.e., $\delta\mathcal{C}=k_{d}S$. Conversely, when $d=2$ the complexity was found to be a fixed constant, independent of the temperature.
@@ 7.4 Generalizations of the holographic proposals
We focus on three generalizations: (1) the extension to the case of subregions, (2) settings with defects or boundaries, and (3) applications to de Sitter spacetime. It is harder to define and test ideas about holographic complexity in these settings than in the standard ones discussed in the previous subsections. Nevertheless, they add information about which notion of quantum complexity should be expected on the field theory side of the holographic duality. [^34] Field theory studies of complexity related to these three extensions appear in, e.g., Agon:2018zso, Caceres:2019pgf for subregions, and Chapman:2018bqj for defects. See also Czech:2016nxc for a relevant tensor network construction and Lin:2022nss for quantum duals to de Sitter spacetime. These constructions usually involve simplified models such as counting arguments for random circuits or even Gaussian free field theories and so can only be compared to the holographic results in a qualitative way.
@@
Consider a mixed state associated with a subregion $\mathcal{A}$ of the boundary. The CV proposal can be extended to this case by computing the maximal volume of a codimension-one surface delimited by the subregion $\mathcal{A}$ and its HRT surface Alishahiha:2015rta. According to the bulk reconstruction paradigm (see Harlow:2018fse for a review), the information carried by the reduced density matrix $\rho_{\mathcal{A}}$ of a mixed state is holographically encoded by the entanglement wedge (EW), which is the bulk domain of dependence of a region that is enclosed within $\mathcal{A}$ and its HRT surface Headrick:2014cta. Therefore, in the case of CA (CV2.0) conjecture, subregion complexity was defined as the gravitational action (spacetime volume) in the bulk region given by the intersection of the WDW patch with the Entanglement Wedge Carmi:2016wjl.
@@
The general structure of the CV and CA proposals has been classified for a ball-shaped subregion $\mathcal{B}$ in arbitrary dimensions, giving a leading UV divergence similar to the full system Carmi:2016wjl. That is, $\mathcal{C}_{\rm sub}\propto k_{d}C_{T}\mathcal{V}(\mathcal{B})/\delta^{d-1},$ where $k_{d}$, the same constant as in eq. (125), carries information on various ambiguities and the dimensionality of the system, $C_{T}$ is the central charge, and $\mathcal{V}(\mathcal{B})$ is the volume of the ball. In general, the subregion complexity proposals have been shown to be superadditive in a pure state; [^35] Consider any holographic complexity conjecture $\mathcal{C}(\Sigma)$ defined on a spacetime region anchored to a boundary Cauchy slice $\Sigma$. Let us split the Cauchy slice $\Sigma$ into a subregion $\mathcal{A}$ and its complement $\bar{\mathcal{A}}$. Denote with $\mathcal{C}(\mathcal{A})$ and $\mathcal{C}(\bar{\mathcal{A}})$ the corresponding subregion complexities, we say that holographic complexity is superadditive if $\mathcal{C}(\Sigma)>\mathcal{C}(\mathcal{A})+\mathcal{C}(\bar{\mathcal{A}})\,$, (126) and subadditive if the inequality is inverted. CA is an exception, since one can tune the counter term scale $L_{\rm{ct}}$ for the null-boundaries (defined below equation (114)) to either get a subadditive or superadditive behaviour Caceres:2019pgf, Agon:2018zso, Alishahiha:2018lfv, Caceres:2018blh, Auzzi:2019fnp. However, restricting the counter term scale $L_{\rm{ct}}$ such that the leading divergence in CA is positive, results in CA being superadditive too. Subsequent studies of holographic complexity associated with subregions were carried out, e.g., in Alishahiha:2015rta, Carmi:2016wjl, Ben-Ami:2016qex, Abt:2017pmf, Roy:2017kha, Roy:2017uar, Bakhshaei:2017qud, Agon:2018zso, Alishahiha:2018lfv, Caceres:2018luq, Abt:2018ywl, Chen:2018mcc, Bhattacharya:2019zkb, Caceres:2019pgf, Auzzi:2019fnp, Auzzi:2019vyh, Auzzi:2019mah, Sato:2021ftf. Relatedly, proposals for a microscopic dual interpretation were investigated, e.g., in Refs. Agon:2018zso, Chapman:2018hou, Balasubramanian:2018hsu, Caceres:2019pgf, Chapman:2019clq, Camargo:2018eof.
@@
The proposed formulae for subregion complexity have an interesting behavior in the BTZ black hole geometry. Given a boundary subregion composed by $q$ segments of total length $\ell_{\rm tot}$, the CV proposal reads Abt:2017pmf
@@
where $c$ is the central charge, $\chi$ the Euler characteristic of the maximal surface, and $\delta$ a UV cutoff regulator. Interestingly, this formula is topological, i.e., it does not depend on the temperature. This feature is not shared by subregion CV in higher dimensions Alishahiha:2015rta, Ben-Ami:2016qex.
The subregion CA and CV2.0 conjectures for a single segment of length $\ell$ in the BTZ background were studied in Auzzi:2019vyh
@@
where $L_{\rm ct}$ is a length scale introduced by a counterterm on null surfaces in the gravitational action, see the discussion below Eq. (114), and $S_{\rm EE}$ is the entanglement entropy of the segment. These results present an appealing feature in that the complexity contains a piece proportional to the entanglement entropy plus an additional constant which could naively depend on the topology as in the CV conjecture results (127). However, an analysis of CA and CV2.0 conjectures in the case of multiple segments in the BTZ background reveals an intricate dependence on the subregion sizes and temperature (see Eqs. (4.23)–(4.24) of Auzzi:2019vyh), which does not follow the same pattern for the dependence of complexity on the entanglement entropies as the one in Eqs. (128a)–(128b) Auzzi:2019vyh, Camargo:2020yfv. Moreover, we do not expect a simple linear relation between complexity and entanglement entropy in dynamical situations, since, as we argued at the beginning of Sec. 7 the two quantities evolve very differently with time.
@@
The first is the two-sided Randall-Sundrum (RS) model, a gravity solution where two patches of AdS spacetime are glued together along a thin codimension-one brane Chapman:2018bqj. The second is the AdS/BCFT model, where an end-of-the-world brane, corresponding to a boundary in the dual CFT, delimits vacuum AdS Braccia:2019xxi, Sato:2019kik. The last is Janus AdS, a dilatonic deformation of AdS spacetime dual to an interface CFT and arising from the dimensional reduction of a solution in type IIB supergravity Auzzi:2021nrj, Auzzi:2021ozb.
@@
22-sided Randall-Sundrum $\frac{2}{3}c\,\eta_{\rm RS}\,\log\left(\frac{\ell}{\,\delta}\right)+\text{finite}$ 0
@@
Table 1: Subregion complexity of formation $\Delta\mathcal{C}$ for an interval of length $\ell$ in the cases of CV and CA, compared to empty AdS spacetime. The coefficients of the log divergences $\eta,\tilde{\eta}$ depend on the details of the defect or boundary, see Chapman:2018bqj, Braccia:2019xxi, Sato:2019kik, Auzzi:2021nrj, Auzzi:2021ozb.
Holographic complexity was studied in Janus AdS geometry under different regularization schemes Baiguera:2021cba. In dimensions $d=2,4$, the authors found a scheme-independent logarithmic divergence, while in odd-dimensions a scheme-independent finite term was observed instead. This result parallels the behavior of the entanglement entropy of smooth subregions, see e.g., Casini2009RevFreeEnt.
@@
De Sitter spacetime is a maximally symmetric solution of Einstein’s equations with a positive cosmological constant. It describes the early stages of evolution of our universe, and possibly the late stages if there is a positive cosmological constant (e.g., see the reviews Spradlin:2001pw, Anninos:2012qw, Galante:2023uyf). Since there is no timelike boundary in de Sitter spacetime, it is challenging to find a holographic interpretation of this geometry in terms of a microscopic quantum mechanical system. Nevertheless, the physical relevance of this spacetime stimulated several attempts Gibbons:1977mu, Bousso:1999dw, Banks:2000fe, Bousso:2000nf, Banks:2001yp, Strominger:2001pn, Balasubramanian:2001nb, Balasubramanian:2002zh, Banks:2002wr, Leblond:2002ns, Leblond:2002tf, Parikh:2002py, Dyson:2002nt, Dyson:2002pf, Parikh:2004wh, Banks:2005bm, Freivogel:2005qh, Banks:2006rx, Lowe:2010np, Anninos:2011af, Fischetti:2014uxa, Anninos:2017hhn, Anninos:2018svg, Banks:2018ypk, Dong:2018cuv, Lewkowycz:2019xse, Banks:2020zcr, Mirbabayi:2020grb, Anninos:2020cwo, Anninos:2020hfj, Susskind:2021omt, Susskind:2021dfc, Coleman:2021nor, Susskind:2021esx, Shaghoulian:2021cef, Shaghoulian:2022fop, Lin:2022nss, Banihashemi:2022htw, Silverstein:2022dfj, Chandrasekaran:2022cip, Witten:2023xze, Witten:2023qsv, Rahman:2023pgt, Narovlansky:2023lfz, Batra:2024kjl, Verlinde:2024znh, Verlinde:2024zrh, Tietto:2025oxn. Recent developments suggest that the inclusion of an artificial timelike boundary in de Sitter spacetime is needed for the consistency of thermodynamics Banihashemi:2022htw, microstate counting Lewkowycz:2019xse, Coleman:2021nor, Batra:2024kjl, and to take into account the role played by an observer Anninos:2011af, Chandrasekaran:2022cip, Witten:2023xze, Witten:2023qsv, Balasubramanian:2023xyd, Balasubramanian:2021wgd, Tietto:2025oxn. These observations led to a revival of static patch holography, according to which a dual quantum theory lives on a codimension-one timelike surface (the stretched horizon) located just inside the cosmological horizon Bousso:2000nf, Parikh:2004wh, Banks:2006rx, Dong:2018cuv, Shaghoulian:2021cef, Susskind:2021esx, Susskind:2021dfc, Shaghoulian:2022fop, Rahman:2023pgt, Narovlansky:2023lfz, Verlinde:2024znh, Verlinde:2024zrh. The holographic complexity proposals introduced in subsection 7.1 can be applied to de Sitter spacetime by requiring that the geometric objects of interest are anchored at the stretched horizons on the two sides of the geometry Susskind:2021esx. The study of these generalizations is speculative, but may hint towards the expected properties of the quantum mechanical dual to de Sitter space.
@@
A main novelty, compared to the AdS case, is that CV, CV2.0 and CA conjectures in dS space include contributions from spacetime regions close to timelike infinity, and therefore diverge at a finite critical time, e.g., see Eq. (3.17) of Jorstad:2022mls. [^36] One can regularize both holographic complexity and its rate by introducing a cutoff surface at future infinity of dS spacetime. After introducing the cutoff, holographic complexity grows linearly, e.g., see Eq. (3.30) of Jorstad:2022mls. It was argued in Susskind:2021esx that this behavior should be interpreted as a hyperfast growth of complexity corresponding to circuits where the gates involve a large number of qubits at each step in the time evolution (e.g., see Lin:2022nss). The hyperfast growth also occurs in two dimensions Anegawa:2023wrk and in models of inflation where a bubble of de Sitter spacetime is inside an AdS geometry Auzzi:2023qbm. [^37] Note however, that in two dimensions, the CV proposal yields a diverging rate of change in complexity as the critical time is approached, but the complexity itself remains finite in this limit. [^38] Holographic complexity has also been studied in other cosmological models, for instance the case of Kasner-like singularities Narayan:2024fcp. However, there are two significant exceptions to this trend. First, there is a class of codimension-one CAny observables that exhibits linear or exponential growth persisting forever, without a divergent rate at finite time Aguilar-Gutierrez:2023zqm. Furthermore, in any dimension of empty de Sitter space, the volume observable relevant to the CV conjecture does not strictly speaking “exist” if we anchor the relevant surfaces on the observer’s world line. To address this, we can instead anchor the surface to a “stretched horizon” at some radius $r_{\text{st}}$. But in this case the volume stops “existing” at a time $t_{\rm crit}$ proportional to $L\,\mathrm{arctanh}(r_{\text{st}}/L)$, where $L$ is the de Sitter scale. So, in addition to the dependence on $L$ and therefore on temperature, the critical time also depends on the radius of the “stretched horizon” that regularizes the problem. If the stretched horizon is very close to the de Sitter horizon, it can be made arbitrarily large Jorstad:2022mls. Similarly, the CV conjecture is strictly speaking only defined for a short time of the order of the inverse temperature in the case of two-dimensional centaur geometries, i.e., gravitational models where de Sitter spacetime is glued to an asymptotic AdS region with a standard timelike boundary Chapman:2021eyy, because the maximal volume surfaces no longer exist after the critical time unless they are regularized in some way. It would be interesting to understand whether these behaviors of the CV observables have an interpretable origin in a dual description.